Reggie is 5 years old than Lani. Five years ago the sum of their ages was 25. How old are they now?
Answer by alicealc(293)

Reggie = R
Lani = L

if it means that Reggie is 5 years older than Lani, then:
R = 5 + L
Five years ago:
R - 5 + L - 5 = 25
5 + L - 5 + L - 5 = 25
2L - 5 = 25
2L = 25 + 5
2L = 30
L = 30/2 = 15
R = 5 + 15 = 20

so, now Reggie is 20 years old, Lani is 15 years old